According to the CDC, if one member of the household contracts the
chickenpox, 90% of family member who have never had the disease or a
vaccine will contract the chickenpox. Suppose four members are exposed
for the first time. What is the probability that three or fewer family
members contract the chickenpox?

Answer by stanbon(75887)

According to the CDC, if one member of the household contracts the chickenpox, 90% of family member who have never had the disease or a vaccine will contract the chickenpox. Suppose four members are exposed for the first time. What is the probability that three or fewer family members contract the chickenpox?
----------
It's a binomial problem with n=4, p=0.9, 0<=x<=3
If you have a TI calculator use binomcdf(4,0.9,3)=0.3439
---------
If you do not have a TI calculator use 1-P(4 contract the chickpox)
=1-4C4 (0.9)^4(0.1)^0= 1-0.6561=0.3439
=============
Cheers,
Stan H.